OpenStudy (anonymous):

A juice processing operation uses steam to heat the fruit juice up to a temperature of 96̊C prior to packaging it in drink boxes. They need to heat 3,250 kg of apple juice (10.5% solids) from an initial storage temperature of 39̊F to its heat treatment temperature. How much saturated steam at 617.8 kPa would be needed for this task (to the nearest kg)? In this case, the condensate is to be discharged from the process without any further cooling to recover any of its heat and is returned to the “steam plant” where it is reheated and turned back into steam

OpenStudy (anonymous):

What is the heat capacity of the juice with 10.5% solids?

OpenStudy (anonymous):

3.83525 kJ/Kg C

OpenStudy (anonymous):

Lemme see. It would take 3.83525 kj/kg K x3250 kg x 57 C = 710480.063 kJ of heat energy. So...

OpenStudy (anonymous):

Please check the heat capacity of steam at 617.8kPa

OpenStudy (anonymous):

for that I have h= hv - hc h = 2758.1 - 675.55 = 2 082.55

OpenStudy (anonymous):

did i do that correctly?

OpenStudy (anonymous):

I'll be honest and just say that I don't remember how to calculate all that from a steam table. It's been a long time and I'll have to refresh my memory Sorry

OpenStudy (anonymous):

that's ok thanks